OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. A(4, 1), B(5, 2)

OpenStudy (anonymous):

use this eq \[\frac{ y-y1 }{ x-x1 }=\frac{ y2-y1 }{ x2 -x1}\] here inn your case y1=1,y2=2,x1=4 and x2=5

OpenStudy (anonymous):

so just plug it in?

OpenStudy (anonymous):

yes...and the solve it simply by cross multiplying...

OpenStudy (anonymous):

can u show me just this one please because i have like 16 more

OpenStudy (anonymous):

yes y not...just a wait...

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (anonymous):

ill be here wenever your ready

OpenStudy (anonymous):

aree you there.?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

here is the sol..

OpenStudy (anonymous):

yyou get it,,,

OpenStudy (anonymous):

kinda!!!!

OpenStudy (anonymous):

kinda mean..?

OpenStudy (anonymous):

am not american..but a pakisttani boy...

OpenStudy (anonymous):

i mean i dont get the cross multiplying!!! wait wat did i get for a final answer its kinda blury

OpenStudy (anonymous):

okay..